# SuperBPE: Space Travel for Language Models

### v1

@@ Abstract
**footnotetext: Equal contribution.The assumption across nearly all language model (LM) tokenization schemes is that tokens should be subwords, i.e., contained within word boundaries. While providing a seemingly reasonable inductive bias, is this common practice limiting the potential of modern LMs? Whitespace is not a reliable delimiter of meaning, as evidenced by multi-word expressions (e.g., by the way), crosslingual variation in the number of words needed to express a concept (e.g., spacesuit helmet in German is raumanzughelm), and languages that do not use whitespace at all (e.g., Chinese). To explore the potential of tokenization beyond subwords, we introduce a “superword” tokenizer, SuperBPE, which incorporates a simple pretokenization curriculum into the byte-pair encoding (BPE) algorithm to first learn subwords, then superwords that bridge whitespace. This brings dramatic improvements in encoding efficiency: when fixing the vocabulary size to 200k, SuperBPE encodes a fixed piece of text with up to 33% fewer tokens than BPE on average. In experiments, we pretrain 8B transformer LMs from scratch while fixing the model size, vocabulary size, and train compute, varying only the algorithm for learning the vocabulary. Our model trained with SuperBPE achieves an average +4.0% absolute improvement over the BPE baseline across 30 downstream tasks (including +8.2% on MMLU), while simultaneously requiring 27% less compute at inference time. In analysis, we find that SuperBPE results in segmentations of text that are more uniform in per-token difficulty. Qualitatively, this may be because SuperBPE tokens often capture common multi-word expressions that function semantically as a single unit. SuperBPE is a straightforward, local modification to tokenization that improves both encoding efficiency and downstream performance, yielding better language models overall.111Code and artifacts will become available at https://superbpe.github.io/.
@@ 1 Introduction
While seemingly reasonable, is this common practice a good one? Whitespace is an unreliable delimiter of meaning—many groups of words (e.g., a lot of or search engine) function semantically as single units, and English speakers store thousands of such multi-word expressions in their mental lexicon (Church, 2011; Contreras Kallens & Christiansen, 2022). Crosslingually, there is considerable variation in whether a given meaning is conveyed by a single word or several words. At the extreme, languages such as Chinese and Japanese do not use whitespace at all, so that tokens in these languages can span multiple words or even entire sentences (e.g., in the tokenizers of GPT-4o [OpenAI, 2024] or DeepSeekv3 [DeepSeek-AI et al., 2025]), yet this has seemingly not hindered LMs from performing well on these languages. Including multi-word tokens promises to be beneficial in several ways: it can lead to shorter token sequences, lowering the computational costs of LM training and inference, and may also offer representational advantages by segmenting text into more semantically cohesive units (Salehi et al., 2015; Otani et al., 2020; Hofmann et al., 2021).
@@ 2.1 Background on BPE
After pretokenization,the iterative learning algorithm begins. T the training text is first split into bytes; the starting vocabulary is the set of all bytes. Then, the frequencies of all pairs of neighboring tokens are recorded, and the most frequent pair is merged into a single, new token at every position in the text where it occurs. The newly merged token is added to the vocabulary. For instance, if the merge is (t, he), then all instances of the token sequence [t, he] will be replaced with the, which is added to the vocabulary. The token pair frequencies are then updated, and the next most frequent pair is again merged into a new token. This continues until vocabulary reaches the target size .
@@ 2.2 SuperBPE tokenization
SuperBPE introduces a simple intervention on the pretokenization step, separating tokenizer training into two discrete phases so that the tokenizer first learns subwords (by using pretokenization to prevent merges across whitespace) and then superwords (by lifting this restriction). Stage 1 is equivalent to regular BPE training and continues up to a certain vocabulary size , which we call the transition point (). In stage 2, tokenizer training resumes from the vocabulary learned so far, but this time the pretokenization step is skipped. As a result, token pairs bridging whitespace are considered, making it possible for superwords to be added to the vocabulary. Intuitively, we wish for our tokenizer to first learn base units of semantic meaning, then combine these units into common sequences for a much more efficient vocabulary. Note that corresponds to BPE, and corresponds to a naive revision of BPE that foregoes whitespace pretokenization at any point in training.
@@ 3.1 Setup
We first pretrain 8B models with BPE and SuperBPE tokenizers. We borrow the exact Olmo2 7B training configuration,333Olmo2 7B has 7.30B parameters while our 8B BPE and SuperBPE models have 8.12B parameters due to their increased vocabulary size. including the model architecture, training hyperparameters, and the pretraining corpus, but reduce the total number of training steps to correspond to 330B tokens (compared to 4T) for the sake of compute. Following prior work (Pagnoni et al., 2024), we also fix the effective context size (measured in bytes) that each model sees. This corrects for any advantage that a model might have by seeing more textual context for the same next-token prediction. As more efficient models will have a shorter context length in tokens, the training steps are adjusted accordingly to match the total train FLOPs at the end of training.444In practice, models using our more efficient tokenizers could shift some or all of the “saved” context FLOPs to longer effective contexts instead of more training steps. Note that in this setting, a SuperBPE model uses fewer inference FLOPs compared to a same-sized BPE model, due to its higher encoding efficiency.
@@ 4.1 Language modeling
Following previous work (Biderman et al., 2023; Xue et al., 2022; Yu et al., 2023; Wang et al., 2024), we evaluate language modeling performance using bits-per-byte (BPB), which normalizes the loss by the encoding efficiency of the tokenizer to fairly compare models with different tokenizers. This is necessary because longer tokens, on average, contain more information and therefore are more difficult to predict. Bits-per-byte is defined as where is the length of the text in bytes and is the sum of the cross-entropy loss over the entire text.777Bits-per-byte of different models are considered comparable because total cross-entropy loss is a universal quantity representing the number of additional bits required to reconstruct the text given the model. This quantity is normalized by the number of bytes for easier interpretation. We find that BPE 8B, SuperBPE 8B ( 180k), and SuperBPE 11B attain 0.7465, 0.7482, and 0.7445 BPB, respectively, at the end of training. Though the numbers are all close, the rankings of the models according to BPB and downstream task performance are not consistent.
@@ 4.2 Fine-grained analysis of bits-per-byte
Why does the SuperBPE 8B model achieve slightly higher normalized language modeling loss (§4.1) than the baseline BPE model, despite outperforming it on a wide variety of downstream tasks (§3.2)? To investigate this, we plot the distribution of per-token BPB for both models on data sampled from the pretraining data mixture in Figure 4. Tokens with BPB were predicted correctly by the model with high confidence, while those with high BPB mean the model placed low probability on the correct next-token.
@@
We find that the reduction in low-loss tokens is attributable to a small set of extremely common words that the BPE model can easily predict, but are not available to SuperBPE as they are merged into larger superword tokens. For instance, the tokens _the, _of, and _to (the three most common words in the corpus) appear an order of magnitude more often under BPE than SuperBPE in the same corpus of text. When excluding these three terms alone, the BPB ranking reverses, with SuperBPE achieving 0.02 lower BPB than BPE.
@@ 4.3 Scaling
To characterize the scaling behavior of SuperBPE, we also perform experiments at smaller scales.888For scaling, we focus on BPB, since our downstream evals are too noisy for our small models to make meaningful comparisons. We train baseline models at 680M and 1.9B, and scale the base number of training tokens proportionately to the number of parameters. We also perform runs at , , and the base number of tokens to observe the trend with respect to training duration. Then, we train two SuperBPE models that match the training budget of each baseline BPE model: one that matches the baseline in parameter count (analogous to SuperBPE 8B), and a larger model that matches in both train and inference compute (analogous to SuperBPE 11B). We focus on the 180k tokenizer to reduce complexity.
@@ 5 Related Work
Prior work has explored processing text at multiple levels of granularity (Lai et al., 2021; Zhang et al., 2021) or creating multi-word tokens through frequency-based identification of -grams (Gee et al., 2023; Kumar & Thawani, 2022). However, these were explored in limited experimental contexts (mainly for machine translation) and had mixed effectiveness. Dagan et al. (2024) find that naively skipping pretokenization severely hurts model performance. In concurrent work, Huang et al. (2025) disentangle the input and output vocabulary, expanding only the input vocabulary to include multi-gram tokens. Their method requires significant modifications of the LM input component and considers fixed length of multi-gram tokens.
@@
Multi-token prediction (MTP) equips LMs with some extra parameters to predict multiple tokens in a single time step (Qi et al., 2020; Gloeckle et al., 2024). This idea was recently adopted by Deepseek-V3 to encourage planning ahead, though the MTP module is discarded at inference-time. The effectiveness corroborates that LMs are capable of predicting more than one subword in a forward pass. However, these approaches fix the number of tokens predicted in each time step and require modifications to the architecture and training objective. The benefits of MTP and superword tokens may be orthogonal, and we see this as a promising avenue for future research.
@@ Acknowledgments
We would like to thank Alex Fang for pretraining advice, Vivek Ramanujan for helping debug our distributed training setup, Ian Magnusson for helpful comments on LM evaluation, and Zhaofeng Wu, Alexander Fang, and Xiaochuang Han for feedback on drafts. We would also like to thank Luca Soldaini, Gonçalo Faria, Shrimai Prabhumoye, Matt Jordan, Artidoro Pagnoni, Mike Lewis, Shannon Shen, and the UW NLP community for valuable conversations about this work. Both co-first authors, AL and JH, are supported by the NSF Graduate Research Fellowship Program. JH and SO are supported in part by Microsoft Grant for Customer Experience Innovation. This work was supported in part by NSF DMS-2134012, ONR N00014-24-1-2207, and NVIDIA resources provided through the National AI Research Resource Pilot (NAIRR).
@@ A.3 Evaluation Suite
All evaluations are based on decoding from the model and scoring the generation by either comparing it to the ground truth or evaluating its functional correctness (in the case of coding tasks). For multiple choice (MC) tasks, we observe that essentially 100% of model generations are valid answer choices, especially for later checkpoints. We choose not to use model-assigned probabilities to assess correctness partly due to surface form competition (Holtzman et al., 2021), where the same answer may be expressed in different forms (USA versus United States), and partly because it is less reflective of the true generation quality. For open-ended tasks, we check whether the generated output contains an exact match to the ground truth answer (EM), and for coding tasks, we report pass@10.
@@
contains simple arithmetic problems (Brown et al., 2020).999https://huggingface.co/datasets/EleutherAI/arithmetic We use the 2da, 2dm, and 2ds splits for addition, multiplication, and division of (up to) 2-digit numbers.
@@
contains open-ended questions from the “Jeopardy!” quiz show. 101010https://www.kaggle.com/datasets/tunguz/200000-jeopardy-questions
@@
contain instructions that ask the model to produce a particular string (Austin et al., 2021).
@@ C.3 Distributional Distortion at the Prompt Boundary
Several fixes for this issue have been proposed, including token healing (Lundberg, 2023), correction procedures at decoding time (Phan et al., 2024), and enumeration of all relevant segmentations of the prompt (Vieira et al., 2024). We leave a detailed comparison of these techniques to future work. Additionally, the issue does not apply at all to models that separate the user’s input from the model’s response using special tokens, as is typical for chat models.

### v2

@@ Abstract
**footnotetext: Equal contribution.The assumption across nearly all language model (LM) tokenization schemes is that tokens should be subwords, i.e., contained within word boundaries. While providing a seemingly reasonable inductive bias, is this common practice limiting the potential of modern LMs? Whitespace is not a reliable delimiter of meaning, as evidenced by multi-word expressions (e.g., by the way), crosslingual variation in the number of words needed to express a concept (e.g., spacesuit helmet in German is Raumanzughelm), and languages that do not use whitespace at all (e.g., Chinese). To explore the potential of tokenization beyond subwords, we introduce a “superword” tokenizer, SuperBPE, which incorporates a simple pretokenization curriculum into the byte-pair encoding (BPE) algorithm to first learn subwords, then superwords that bridge whitespace. This brings dramatic improvements in encoding efficiency: when fixing the vocabulary size to 200k, SuperBPE encodes a fixed piece of text with up to 33% fewer tokens than BPE on average. In experiments, we pretrain 8B transformer LMs from scratch while fixing the model size, vocabulary size, and train compute, varying only the algorithm for learning the vocabulary. Our model trained with SuperBPE achieves an average +4.0% absolute improvement over the BPE baseline across 30 downstream tasks (including +8.2% on MMLU), while simultaneously requiring 27% less compute at inference time. In analysis, we find that SuperBPE results in segmentations of text that are more uniform in per-token difficulty. Qualitatively, this may be because SuperBPE tokens often capture common multi-word expressions that function semantically as a single unit. SuperBPE is a straightforward, local modification to tokenization that improves both encoding efficiency and downstream performance, yielding better language models overall.111Code and artifacts are available at https://superbpe.github.io/.
@@ 1 Introduction
While seemingly reasonable, is this common practice a good one? Whitespace is an unreliable delimiter of meaning (Martin, 2017)—many groups of words (e.g., a lot of or search engine) function semantically as single units, and English speakers store thousands of such multi-word expressions in their mental lexicon (Church, 2011; Contreras Kallens & Christiansen, 2022). Crosslingually, there is considerable variation in whether a given meaning is conveyed by a single word or several words. At the extreme, languages such as Chinese and Japanese do not use whitespace at all, so that tokens in these languages can span multiple words or even entire sentences (e.g., in the tokenizers of Gpt-4o [OpenAI, 2024] or DeepSeekv3 [DeepSeek-AI et al., 2025]), yet this has seemingly not hindered LMs from performing well on these languages. Including multi-word tokens promises to be beneficial in several ways: it can lead to shorter token sequences, lowering the computational costs of LM training and inference, and may also offer representational advantages by segmenting text into more semantically cohesive units (Salehi et al., 2015; Otani et al., 2020; Hofmann et al., 2021).
@@ 2.1 Background on BPE
After pretokenization, the iterative learning algorithm begins. The training text is first split into bytes; the starting vocabulary is the set of all bytes. Then, the frequencies of all pairs of neighboring tokens are recorded, and the most frequent pair is merged into a single, new token at every position in the text where it occurs. The newly merged token is added to the vocabulary. For instance, if the merge is (t, he), then all instances of the token sequence [t, he] will be replaced with the, which is added to the vocabulary. The token pair frequencies are then updated, and the next most frequent pair is again merged into a new token. This continues until vocabulary reaches the target size .
@@ 2.2 SuperBPE tokenization
SuperBPE introduces a simple intervention on the pretokenization step, separating tokenizer training into two discrete phases so that the tokenizer first learns subwords (by using pretokenization to prevent merges across whitespace) and then superwords (by lifting this restriction). Stage 1 is equivalent to regular BPE training and continues up to a certain vocabulary size , which we call the transition point (). In stage 2, tokenizer training resumes from the vocabulary learned so far, but this time whitespace pretokenization is skipped. As a result, token pairs bridging whitespace are considered, making it possible for superwords to be added to the vocabulary. Intuitively, we wish for our tokenizer to first learn base units of semantic meaning, then combine these units into common sequences for a much more efficient vocabulary. Note that corresponds to BPE, and corresponds to a naive revision of BPE that foregoes whitespace pretokenization at any point in training.
@@ 3.1 Setup
We first pretrain 8B models with BPE and SuperBPE tokenizers. We use the Olmo2 7B (OLMo et al., 2024) training configuration,333Olmo2 7B has 7.30B parameters while our 8B BPE and SuperBPE models have 8.12B parameters due to their increased vocabulary size. including the model architecture, training hyperparameters, and pretraining corpus, but reduce the total number of training steps to correspond to 330B tokens (compared to 4T) for the sake of compute. Following prior work (Pagnoni et al., 2024), we also fix the effective context size (measured in bytes) for each model. This prevents SuperBPE models from gaining an advantage by seeing more textual context for the same next-token prediction (Xiong et al., 2024). As more efficient models will have a shorter context length in tokens, the training steps are adjusted accordingly to match the total train FLOPs at the end of training.444In practice, models using our more efficient tokenizers could shift some or all of the “saved” context FLOPs to longer effective contexts instead of more training steps. Note that in this setting, a same-sized SuperBPE model uses fewer inference FLOPs compared to the BPE model.
@@ 4.1 Language modeling
Following prior work (Biderman et al., 2023; Xue et al., 2022; Yu et al., 2023; Wang et al., 2024), we evaluate language modeling performance using bits-per-byte (BPB), which normalizes the loss by the encoding efficiency of the tokenizer to fairly compare models with different tokenizers. This is necessary because longer tokens, on average, contain more information and therefore are more difficult to predict. Bits-per-byte is defined as where is the length of the text in bytes and is the sum of the cross-entropy loss over the entire text.777Bits-per-byte of different models are considered comparable because total cross-entropy loss is a universal quantity representing the number of additional bits required to reconstruct the text given the model. This quantity is normalized by the number of bytes for easier interpretation. We find that BPE 8B, SuperBPE 8B ( 180k), and SuperBPE 11B attain 0.7465, 0.7482, and 0.7445 BPB, respectively, at the end of training. Though the numbers are all close, the rankings of the models according to BPB and downstream task performance are not consistent.
@@ 4.2 Fine-grained analysis of bits-per-byte
Why does the SuperBPE 8B model achieve slightly higher normalized language modeling loss (§4.1) than the baseline BPE model, despite outperforming it on a wide variety of downstream tasks (§3.2)? To investigate this, we plot the distribution of per-token BPB888The per-token BPB is the per-token loss (in bits) divided by the average encoding efficiency. for both models on data sampled from the pretraining data mixture in Figure 4.
@@
We find that the reduction in low-loss tokens is attributable to a small set of extremely common words that the BPE model can easily predict, but are not available to SuperBPE as they are merged into larger superword tokens. For instance, the tokens _the, _of, and _to (the three most common words in the corpus) appear an order of magnitude more often under BPE than SuperBPE in the same corpus of text. When excluding these three token types alone, the BPB ranking reverses, with SuperBPE achieving 0.02 lower BPB than BPE.
@@ 4.3 Scaling
To characterize the scaling behavior of SuperBPE, we also perform experiments at smaller scales.999For scaling, we focus on BPB, since our downstream evals are too noisy for our small models to make meaningful comparisons. We train baseline models at 680M and 1.9B, and scale the base number of training tokens proportionately to the number of parameters. We also perform runs at , , and the base number of tokens to observe the trend with respect to training duration. Then, we train two SuperBPE models that match the training budget of each baseline BPE model: one that matches the baseline in parameter count (analogous to SuperBPE 8B), and a larger model that matches in both train and inference compute (analogous to SuperBPE 11B). We focus on the 180k tokenizer to reduce complexity.
@@ 5 Related Work
Prior work has explored processing text at multiple levels of granularity (Lai et al., 2021; Zhang et al., 2021) or creating multi-word tokens through frequency-based identification of -grams (Gee et al., 2023; Kumar & Thawani, 2022). However, these were explored in limited experimental contexts (mainly for machine translation) and had mixed effectiveness. Naively disabling pretokenization in BPE has been found to severely hurt model performance (Dagan et al., 2024; Schmidt et al., 2024; Kudo, 2018), though it may be more promising for unigram tokenization (Kudo & Richardson, 2018), as adopted by Jurassic (Lieber et al., 2021) and BloombergGpt (Wu et al., 2023). In concurrent work, Huang et al. (2025) disentangle the input and output vocabulary, expanding only the former to include -gram tokens. Their method requires significant modifications of the LM input component and considers fixed length of -gram tokens.
@@
Multi-token prediction (MTP) equips LMs with some extra parameters to predict multiple tokens in a single time step (Qi et al., 2020; Gloeckle et al., 2024), and was recently adopted by Deepseek-V3, though the MTP module is discarded at inference-time. The effectiveness corroborates that LMs are capable of predicting more than one subword in a forward pass. However, these approaches fix the number of tokens predicted in each time step and require modifications to the architecture and training objective. Nonetheless, the benefits of MTP and superword tokens may be orthogonal.
@@ Acknowledgments
We would like to thank Alex Fang for pretraining advice, Vivek Ramanujan for helping debug our distributed training setup, Ian Magnusson for helpful comments on LM evaluation, and Zhaofeng Wu, Alexander Fang, and Xiaochuang Han for feedback on drafts. We would also like to thank Luca Soldaini, Gonçalo Faria, Shrimai Prabhumoye, Matt Jordan, Artidoro Pagnoni, Mike Lewis, Doug Downey, Shannon Shen, and the UW NLP community for valuable conversations about this work. Both co-first authors, AL and JH, are supported by the NSF Graduate Research Fellowship Program. JH and SO are supported in part by the Microsoft Grant for Customer Experience Innovation. This work was further supported in part by NSF DMS-2134012, NSF CCF-2019844, ONR N00014-24-1-2207, NSF 2113530, and NVIDIA resources provided through the National AI Research Resource Pilot (NAIRR).
@@ A.3 Evaluation Suite
All evaluations are based on decoding from the model and scoring the generation by either comparing it to the ground truth or evaluating its functional correctness (in the case of coding tasks). For multiple choice (MC) tasks, we check whether the predicted answer choice is an exact match (EM) to the target (we observe that effectively 100% of model generations are valid answer choices, especially for later checkpoints). For open-ended tasks, we check whether the generated output contains the ground truth answer exactly, and for coding tasks, we report pass@10.
@@
contains simple arithmetic problems (Brown et al., 2020).101010https://huggingface.co/datasets/EleutherAI/arithmetic We use the 2da, 2dm, and 2ds splits for addition, multiplication, and division of (up to) 2-digit numbers.
@@
contains open-ended questions from the “Jeopardy!” quiz show. 111111https://www.kaggle.com/datasets/tunguz/200000-jeopardy-questions
@@
contains instructions that ask the model to produce a particular string (Austin et al., 2021).
@@ C.3 Distributional Distortion at the Prompt Boundary
Several fixes for this issue have been proposed: at training time, token merges can be randomly dropped (Provilkov et al., 2020; Sims et al., 2025; DeepSeek-AI et al., 2025) to expose LMs to the internal makeup of tokens; at inference time, options include token healing (Lundberg, 2023), algorithmic correction (Phan et al., 2024), and enumeration of all relevant segmentations of the prompt (Vieira et al., 2024). We leave a detailed comparison of these techniques to future work.
Additionally, the issue does not apply at all to models that separate the user’s input from the model’s response using special tokens, as is typical for chat models.
Appendix D Other Related Work
Please see Mielke et al. (2021) for a survey of subword tokenization.
Pretokenization
Pretokenization defines how the text is split in order to prevent certain pairs of tokens from being merged. Gpt-2 (Radford et al., 2019) introduced a regular expression (regex) which defines the pretokenization pattern. These regex strings have gained complexity over time; Gpt-3.5 limits the number of digits in numerical tokens to 3, and allows single punctuation to be merged with the start of words (presumably to accommodate code, as it allows .get to be a single token). Prior work has shown that, for instance, digit pretokenization choices (Nogueira et al., 2021; Thawani et al., 2021; Singh & Strouse, 2024) can significantly impact arithmetic performance. It is also likely that pretokenization affects different languages differently (Velayuthan & Sarveswaran, 2025; Ahia et al., 2023), due to natural statistics of the average word length, which acts as an upper bound on encoding efficiency in that language under subword tokenization. Nonetheless, the effectiveness of many pretokenization choices have not been thoroughly studied.
-gram language models
Internal representation of semantic units
Previous work has showed that the early layers of the LM may “aggregate” information over multi-token entities (e.g., [_New, _York]) into the last token’s (e.g., _York) hidden representation (Meng et al., 2022; Kaplan et al., 2025; Lad et al., 2024). This suggests that LMs naturally learn multi-word representations, and segmentating text into more semantically cohesive units at the input level (e.g., having _New_York as a single token) may simplify this process.